# On the modified $J$-equation

### v3

@@ Abstract.
In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16]. We show that the solvability of the modified $J$-equation is equivalent to the coercivity of the modified $J$-functional on compact Kähler manifolds. For smooth projective toric varieties we establish a Nakai–Moishezon type criterion for the existence of solutions, extending the results of Collins–Székelyhidi [CS17]. As a possible application, by combining our result with Delcroix–Jubert [DJ25], we further discuss a numerical sufficient condition for the existence of extremal Kähler metrics on smooth projective toric varieties.
@@ 1. Introduction
Finding canonical metrics has been one of the central themes in Kähler geometry. Especially several equations involving Hamiltonians have been studied extensively in the last few decades (e.g. [BJ25, FO19, Lah19, Li21, Mab03, TZ00, Zhu00]). Hamiltonians of holomorphic vector fields encode the symmetries of the manifold with respect to a given Kähler metric. When studying canonical metrics, such as Kähler–Einstein or constant scalar curvature Kähler (cscK) metrics, it is natural to take these Hamiltonians into account, since the existence and uniqueness of such metrics are often obstructed by holomorphic automorphisms. By introducing Hamiltonians of holomorphic vector fields, one can treat the variational problem even without a canonical metric.
In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16]:
@@
is independent of the choice of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, and the existence of a solution to (1.2) easily leads to $m_{X}>0$ (see Section 2.2). In particular, if $T$ is trivial then the equation (1.2) matches up with the $J$-equation introduced by Donaldson [Don03] and Chen [Che00]. Although the $J$-equation itself is an intriguing equation, it has sometimes been studied in relation to cscK metrics, as it provides the starting point of the continuity method for solving the cscK equation [CC18, CC21a, CC21b]. In the modified case, Li–Shi [LS16] showed that such a solution $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ to (1.2) exists if and only if there exists $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$ such that
$\big(c_{X}+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0,$ (1.5)
where the positivity of (n1,n1)(n-1,n-1)-forms is understood in the sense of [Dem12, Chapter III, Section 1.A]. The solutions to the equation (1.2) are characterized as the critical points of a strictly convex function $J_{v}^{\omega}$ on $\mathcal{H}(X,\widehat{\chi})^{T}$, called the modified $J$-functional. On the other hand, the modified $K$-energy $K_{v_{\rm ext}}$ is a convex functional on the space of Kähler potentials whose critical points are extremal Kähler metrics. Motivated by [LS16], Delcroix–Jubert [DJ25] recently considered a modified version of the Chen–Tian formula, relating these functionals, and proved that the existence of a solution to the modified $J$-equation implies the $T^{\mathbb{C}}$-coercivity of the modified $K$-energy.
For later arguments, it is convenient to extend (1.2) to a more general equation of the form:
@@
In most cases, we assume that $b\geqslant 0$, but sometimes allow $b$ to be slightly negative. First, we extends the result [LS16, Theorem 3.3] to the generalized equation (1.6) as follows:
@@
In the study of canonical Kähler metrics, a fundamental question is whether the solvability of an equation can be related to the properness or coercivity of a certain energy functional. For instance, Tian’s properness conjecture [Tia94] predicted that the existence of Kähler–Einstein metrics is equivalent to the properness of the $K$-energy. When the manifold has trivial automorphism group, Ding–Tian [DT92] proved one direction (properness $\Rightarrow$ existence), and Tian [Tia97a] confirms a converse direction (existence $\Rightarrow$ properness). In general case, Darvas–Rubinstein [DR17] proposed a refined version of Tian’s properness conjecture and proved it by exploiting a general framework in Finsler geometry. For constant scalar curvature Kähler metrics, a version of Tian’s properness conjecture was proved by Berman–Darvas–Lu [BDL20] and Chen–Cheng [CC18, CC21a, CC21b]. For complex Hessian equations, Collins–Székelyhidi [CS17] obtained a similar result for the $J$-equation, and Chu–Lee [CL21] for the hypercritical deformed Hermitian–Yang–Mills equation.
Motivated by above works, we show the equivalence between the existence and coercivity for the modified $J$-equation:
@@
For the precise definition of $J_{v}^{\omega}$ and coercivity, see Section 2.4. To prove the above theorem, we use the local smoothing and gluing method established in [Che21, CS17]. A new issue arising in the modified case is that even if we approximate the holomorphic vector field $v$ locally by vector fields with constant coefficients, to deal with the approximation error terms, we need a uniform gradient estimate for the potentials along the modified $J$-flow. To avoid this, we consider the product of $X$ with an annulus, add a rotational vector field $w$ to $v$ so as to eliminate the zeroes. We take holomorphic coordinates on which $v+w$ has constant coefficients by using the rectification theorem. Then we perform smoothing, gluing and finally averaging in the rotational direction to cancel out the contribution of $w$ (see Section 4.1.1 for more details).
@@
[…]
@@
When dealing with the contribution of the holomorphic vector field $v$, one need to handle carefully the fact that the normalization of the Hamiltonian (1.1) differs for each subvariety $Y$. Also, when gluing subsolutions, we need to control the derivatives of potential functions along $v$ near each $D_{i}$, using a $T$-equivariant linearization of $\mathcal{O}_{X}(D_{i})$. For this reason, we make the additional assumption that $X$ is projective. As an application, we also obtain:
Theorem 1.7.
Let $X$ be an $n$-dimensional smooth projective toric variety with real torus $T$, and let $\widehat{\chi}$ be a $T$-invariant Kähler form on $X$ with Kähler class $\beta$. Let $\delta_{\beta}^{T^{\mathbb{C}}}$ be the $T^{\mathbb{C}}$-reduced delta invariant of $\beta$ defined by (6.2). Let $v_{\rm ext}$ be the extremal vector field associated with $\beta$ such that $\mathrm{Im}(v_{\rm ext})\in\mathfrak{t}$. Assume that there exists a constant $\varepsilon>0$ satisfying the following properties:
  1. (1)
  2. (2)
    For any $T$-invariant Kähler forms $\chi\in\beta$,
  3. (3)
    For any $p$-dimensional irreducible toric subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ and any $T$-invariant Kähler forms $\chi\in\beta$,
Remark 1.8.
We emphasize that the above theorem is essentially established in [DJ25, Theorem 1.4]. Our contribution is limited in the toric case. Specifically, in the course of the proof, we need to show the existence of a solution to a certain modified $J$-equation, which we reformulate in numerical terms by using Theorem 1.6. Thus, if Conjecture 1.4 is true, Theorem 1.7 still holds for arbitrary compact Kähler manifolds. Also, it should be noted that in the toric case, a stronger statement holds: the conclusions remain valid even if $\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon$ in assumptions (1)–(3) is replaced by $\delta_{\beta}^{T^{\mathbb{C}}}$. This is due to the fact that, in (3), the object under considerations are restricted to toric subvarieties, which form a finite set.
We note that in the case $v=0$, Theorem 1.2, Theorem 1.3 and Theorem 1.6 were previously established in [CS17]. Also, a variant of Theorem 1.7 was obtained in [Che21, Corollary 1.5].
@@
The organization of this paper is as follows. In Section 2 we collect some fundamental properties that will be used throughout this paper. In Section 3 we establish the long-time existence of the modified $J$-flow. In particular we prove the convergence of the flow when $b\geqslant 0$ under the existence of solusolutions, which gives a proof of Theorem 1.1. In Section 4 we show that the existence of solutions is equivalent to the coercivity of the functional $J_{v}^{\omega}$ (Theorem 1.2). As a corollary, we also obtain a proof of Theorem 1.3. In Section 5 we establish a Nakai–Moishezon type criterion for smooth projective toric varieties, which provides the proof of Theorem 1.6. Finally, in Section 6, we give the proof of Theorem 1.7.
@@ 2.3. Subsolutions
In this subsection we study general properties for subsolutions (1.5) related to the regularized maximum and averaging over the torus $T$. We first discuss how the subsolution coindition can be expressed in terms of the operator $P_{\omega}$.
@@
$\chi\geqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j})>0.$
@@
Since $\omega$ is $T$-invariant, the Kähler form $\tau^{\ast}\chi$ can be identified with a positive definite Hermitian matrix by using an $\omega$-orthonormal basis at each point on $X$. So the convexity of $P_{\omega}$ implies that
@@ 3.1. Long-time existence for $b=-\varepsilon<0$
Since $\lambda_{i}$ is uniformly bounded from blow this shows that the right hand side is bounded from above by some uniform constant $C_{4}>0$. Therefore if we set $G:=\log\operatorname{Tr}_{\omega}\chi_{t}-2C_{4}t$ then for any $T\in(0,T_{\max})$ and maximum point $(x_{0},t_{0})$ of $G$ on $X\times[0,T]$ with $t_{0}>0$ we must have
@@ 4.1. Coercivity implies existence
[…]
@@ 4.1.1. Local smoothing
The Kähler current $\chi_{u_{\infty}}$ can be regarded as a solution to (1.2) in a weak sense. Now we establish estimates for local smoothing of $\chi_{u_{\infty}}$ in Euclidean balls. The main difference from the usual $J$-equation lies in the smoothing estimates for the Hamiltonian, which are not straightforward. Even if $v$ is locally approximated by a constant-coefficient vector field, controlling the resulting error term requires a uniform gradient bound $|\nabla u_{k}|_{g}\leqslant C$. While this error term could be eliminated using the rectification theorem, this approach can not be applied directly since $v$ must have a zero on $X$. To address this issue, one may consider a product manifold $\mathcal{X}:=X\times\mathbb{C}$ instead of $X$: let $\omega_{\mathbb{C}}:=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}|\tau|^{2}$ be the standard Kähler form on $\mathbb{C}$ and $w:=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$ a holomorphic vector field on $\mathbb{C}$ whose real part generates the action of the real torus $T^{\prime}\simeq\mathbb{S}^{1}$ on $\mathbb{C}$ by rotations. Set
@@
Since the holomorphic vector field $v+w$ has no zeros on $\mathcal{X}_{3}$, for any fixed $p\in\mathcal{X}_{3}$ we can choose local homomorphic coordinates $(w^{1},\ldots,w^{n+1})$ centered at $p$ such that $\frac{\partial}{\partial w^{1}}=v+w$ by the rectification theorem. In this coordinates, after adjusting the affine term, the local potential function $\phi$ of $\omega_{\mathcal{X}}$ is expressed as $\phi(w)=\overline{w}^{T}Gw+O(|w|^{3})$ for some positive definite Hermitian matrix $G=(G_{i\bar{j}})$. In particular, we have $|v+w|_{g_{\mathcal{X}}}^{2}(p)=G_{1\bar{1}}$, where $g_{\mathcal{X}}$ denotes the Riemannian metric with respect to $\omega_{\mathcal{X}}$. Take $B:=G^{1/2}$ so that the local potential function $\phi$ in the coordinates $z=Bw$ becomes $\phi(z)=|z|^{2}+O(|z|^{3})$. The holomorphic vector field $v+w$ still has constant coefficients in $(z^{1},\ldots,z^{n+1})$, which we denote by $c(v+w)\in\mathbb{C}^{n+1}$. Also a simple computation shows that $|c(v+w)|=|v+w|_{g_{\mathcal{X}}}(p)\leqslant\sup_{\mathcal{X}_{3}}|v+w|_{\omega_{\mathcal{X}}}$.
@@
For any $\varepsilon\in(0,\varepsilon_{7})$ we have
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\leqslant F_{\omega,-\varepsilon}(\chi_{u_{k}})+\varepsilon^{2},$
where the quantity in the left-hand side is calculated on $\mathcal{X}$.

### v4

@@ Abstract.
In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16]. We first show that, on compact Kähler manifolds, the solvability of the modified $J$-equation is equivalent to the coercivity of the modified $J$-functional. Motivated by this characterization, we formulate a Nakai–Moishezon type criterion for the existence of solutions to the modified $J$-equation on general compact Kähler manifolds. We then verify this conjectural criterion in the case of smooth projective toric varieties. This extends the work of Collins–Székelyhidi [CS17] and provides further evidence for the expected algebro-geometric nature of the modified $J$-equation. As a potential application, we combine our results with Delcroix–Jubert [DJ25]. Assuming our conjectural Nakai–Moishezon type criterion holds in general, we obtain a numerical sufficient condition for the existence of extremal Kähler metrics on arbitrary compact Kähler manifolds.
@@ 1. Introduction
1.1. Background and motivation
Finding canonical metrics has been one of the central themes in Kähler geometry. Equations such as the constant scalar curvature Kähler (cscK) equation, the Kähler–Einstein equation, and their various twisted or weighted analogues reveal deep interactions between complex geometry, algebraic stability, and variational principles (e.g. [BJ25, FO19, Lah19, Li21, Mab03, TZ00, Zhu00]). Among these, the $J$-equation, introduced by Donaldson [Don03] and Chen [Che00], plays a crucial role in understanding the relation between energy properness and the existence of cscK metrics. It also serves as an important model in the continuity approach to cscK metrics [CC18, CC21a, CC21b]. An important progress in the analysis of the $J$-equation was achieved by Song–Weinkove [SW08] who proved that the existence of a subsolution is equivalent to the solvability of the equation. Building on this insight, Collins–Székelyhidi [CS17] established a fundamental correspondence between the solvability of the $J$-equation and the coercivity of the associated functional, providing a model for the continuity method in Kähler geometry. A further breakthrough was made by G. Chen [Che21], who established a connection between the solvability of the $J$-equation and algebro-geometric stability. Using Chen’s result [Che21], Song [Son20] obtained a Nakai–Moishezon type numerical criterion for the solvability of the $J$-equation, confirming the Lejmi–Székelyhidi conjecture [LS15].
However, the usual $J$-equation does not incorporate the effect of holomorphic automorphisms, which often obstruct the existence or uniqueness of canonical metrics. To address this, Li–Shi [LS16] introduced the modified $J$-equation :
@@ 1.1. Background and motivation
is independent of the choice of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, and the existence of a solution to (1.2) easily leads to $m_{X}>0$ (see Section 2.2). They proved that the existence of a subsolution $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$:
$\big(c_{X}+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0$ (1.5)
is equivalent to the solvability of the equation (1.2). This equivalence was obtained by studying the modified $J$-flow
$\frac{d}{dt}\phi_{t}=-\operatorname{Tr}_{\chi_{\phi_{t}}}\omega+c_{X}+\theta_{v}^{X}(\chi_{\phi_{t}})$ (1.6)
with initial data $\phi_{0}\in\mathcal{H}(X,\widehat{\chi})^{T}$. The modified $J$-equation is closely related to the study of extremal Kähler metrics, since the latter can be viewed as critical points of the modified $K$-energy. In this sense, the modified $J$-theory may serve as an analytic model for understanding extremal and, more generally, weighted cscK metrics (see Section 1.3).
1.2. Main results
The purpose of this paper is to establish the analytic and numerical foundations of the modified $J$-theory, and to clarify how it extends the previous framework developed by [CS17, LS16]. For later arguments, it is convenient to extend (1.2) to a more general equation of the form:
@@ 1.2. Main results
In most cases, we assume that $b\geqslant 0$, but sometimes allow $b$ to be slightly negative. We extends the result [LS16, Theorem 3.3] to the generalized equation (1.7) as follows:
@@
Our first main result is an analogue of the existence-coercivity correspondence proved for the usual $J$-equation in [CS17].
@@
For the precise definition of $J_{v}^{\omega}$ and coercivity, see Section 2.4. This result shows that the coercivity of the modified $J$-functional completely characterizes solvability, in the same spirit as the coercivity criterion for various canonical metrics (e.g. [BDL20, CC18, CC21a, CC21b, CL21, DR17, DT92, Tia97a]).
@@
Remark 1.6.
In particular, when $T$ is trivial and $n=2$, the modified $J$-equation (1.2) reduces to a complex Monge-Ampère equation:
$(c_{X}\chi_{\phi}-\omega)^{2}=\omega^{2}.$
Consequently, Conjecture 1.4 follows directly from the criterion of Demailly–Păun [DP04] and [Yau78]. In contrast, when $T$ is non-trivial, even for surfaces ($n=2$), the modified $J$-equation (1.2) does not reduce to a complex Monge-Ampère equation, and Conjecture 1.4 remains genuinely nontrivial.
In the case where $T$ is trivial, this was proposed by Lejmi–Székelyhidi [LS15] and subsequently established by Song [Son20], using [Che21] as a key ingredient. We also note that, when the condition (1.10) fails, the formation of singularities of the modified $J$-flow (1.6) in the Calabi ansatz has been studied by Sivaram [Siv24]. Focusing on the toric setting allows for more explicit results. Theorem 1.7 confirms Conjecture 1.4 in this case, providing a concrete demonstration of the expected behavior under these assumptions.
@@
As in [CS17, Theorem 3], the proof proceeds by analyzing the behavior outside the union of toric divisors $D=\cup_{i}D_{i}$, adapting their method to the present context. Unlike [CS17], we need to control the Hamiltonian near each $D_{i}$ during the gluing process, by using a $T$-equivariant linearization of $\mathcal{O}_{X}(D_{i})$. For this reason, we make the additional assumption that $X$ is projective.
1.3. Relation to extremal Kähler metrics
Recall that extremal Kähler metrics are precisely the critical points of the modified $K$-energy $K_{v_{\rm ext}}$, which can be decomposed into the relative entropy and the modified $J$-functional. This decomposition is exactly the Chen–Tian formula [Che00, Tia97a, Tia97b]:
$K_{v_{\rm ext}}=H+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}$.
Using this formula, Li–Shi [LS16] established a sufficient condition related to the modified $J$-equation under which the modified $K$-energy is coercive. Building on this perspective, Delcroix–Jubert [DJ25] further showed that the $T^{\mathbb{C}}$-coercivity of the modified $K$-energy can be reduced to the existence problem for the modified $J$-equation. This reduction is a key step in connecting extremal Kähler metrics with solutions of the modified $J$-equation. Their work, however, addresses a more general setting, namely so-called weighted cscK metrics, which have recently attracted considerable attention in Kähler geometry (e.g. [AJL23, BJ25, HL24, Lah19, NJL25]). It is natural to expect that the theory of the modified $J$-equation can be further adapted in accordance with their framework, potentially yielding applications to the existence problem for such general canonical metrics. We leave a systematic exploration of such general cases to future work. At the same time, the combination of our results with [DJ25] provides a framework for potential applications to extremal Kähler metrics. We explore these applications and related discussions in Section 6.
1.4. Contributions
The main contributions of this paper concern the theory of the modified $J$-equation, with the principal new analytic ingredients being the local smoothing and gluing. Our analysis builds on the techniques developed in [Che21, CS17], which we further adapt to handle the lack of $T$-invariance and the degeneracy of the automorphism flow generated by the holomorphic vector field $v$. These issues form a major obstacle to the construction of subsolutions through local smoothing and gluing procedures. To overcome them, we employ several analytic tools. The first contribution consists of two techniques: the introduction of averaging with respect to the torus action, which restores $T$-invariance, and a careful control of the Hamiltonian near $T$-invariant subvarieties. These together make it possible to carry out the gluing procedures in a way compatible with the $T$-action (cf. Section 2.3 and Theorem 5.2).
The second is the annulus trick, used in the construction of solutions in Theorem 1.2. When establishing the local smoothing estimate, the key is to work in local coordinates that are compatible with the automorphism flow generated by the holomorphic vector field $v$, so that the modified structure is preserved. However, such coordinates can not be defined consistently near the zero set of $v$, where the automorphism action degenerates. To avoid this, we consider the product space $X\times A$, where $A=\{\tau\in\mathbb{C}\mid 1/2<|\tau|<2\}$ is an annulus equipped with a rotational vector field $w=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$. By lifting $v$ to $X\times A$ and adding $w$, we obtain a flow without fixed points. This extended flow allows us to define local coordinates adapted to it, which are then used to perform smoothing and gluing constructions. Combined with an averaging argument along the rotational direction, this yields a subsolution on the slice $X\simeq X\times\{1\}\subset X\times A$. This trick is not merely a computational convenience; it is essential in that it allows us to bypass the highly nontrivial analytic problem of establishing uniform gradient estimates along the modified $J$-flow (see Section 4.1.1 for more details).
1.5. Organization of the paper
@@
The organization of this paper is as follows. In Section 2 we review several basic properties that will be used throughout the paper. Section 3 is devoted to the study of the long-time behavior of the modified $J$-flow; in particular, we prove its convergence when $b\geqslant 0$ under the existence of subsolutions, thereby establishing Theorem 1.1. In Section 4, we show that the existence of solutions is equivalent to the coercivity of the functional $J_{v}^{\omega}$ (Theorem 1.2), from which Theorem 1.3 follows as a corollary. Section 5 develops a Nakai–Moishezon type criterion for smooth projective toric varieties, providing the proof of Theorem 1.7. Finally, in Section 6, we discuss how Conjecture 1.4 can be combined with the results of Delcroix–Jubert [DJ25] to address the existence problem for extremal Kähler metrics.
@@ 2.3. Subsolutions
In this subsection we study general properties for subsolutions (1.5) related to the regularized maximum and averaging over the torus $T$. We first discuss how the subsolution condition can be expressed in terms of the operator $P_{\omega}$.
@@
$\chi\geqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\bigg)>0.$
@@
Since $\omega$ is $T$-invariant, the Kähler form $\tau^{\ast}\chi$ can be identified with a positive definite Hermitian matrix by using an $\omega$-orthonormal basis at each point on $X$. So the convexity of the function $P$ implies that
@@ 3.1. Long-time existence for $b=-\varepsilon<0$
Since $\lambda_{i}$ is uniformly bounded from below this shows that the right hand side is bounded from above by some uniform constant $C_{4}>0$. Therefore if we set $G:=\log\operatorname{Tr}_{\omega}\chi_{t}-2C_{4}t$ then for any $T\in(0,T_{\max})$ and maximum point $(x_{0},t_{0})$ of $G$ on $X\times[0,T]$ with $t_{0}>0$ we must have
@@ 4.1. Coercivity implies existence
for some uniform constant $C>0$ (independent of $\varepsilon$ and $k$). Hence, by choosing a subsequence, we may assume that $u_{k}\to u_{\infty}\in\mathcal{E}^{1}(X,\widehat{\chi})^{T}$ in $L^{1}$ and $\chi_{u_{k}}\to\chi_{u_{\infty}}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty}$ weakly as $k\to\infty$, where $\mathcal{E}^{1}(X,\widehat{\chi})^{T}$ denotes the space of $T$-invariant $\widehat{\chi}-$PSH functions with finite energy (see Remark 2.18 and [BBGZ13, Lemma 3.3]). So by [GZ07, Corollary 1.8] the function $u_{\infty}$ has zero Lelong numbers.
@@ 4.1.1. Local smoothing
The Kähler current $\chi_{u_{\infty}}$ can be regarded as a solution to (1.2) in a weak sense. Now we establish estimates for local smoothing of $\chi_{u_{\infty}}$. Following the general strategy in [Che21, CS17], we locally approximate the Kähler form $\omega$ and the holomorphic vector field $v$ by constant-coefficient ones in each coordinate chart. This approximation makes the smoothing operator commute with differentiation. For the Hessian term, the desired estimate then follows from the monotonicity and convexity of the function $P$. However, the argument for the Hamiltonian term is more delicate. Even if $v$ is locally approximated by a constant-coefficient vector field, controlling the resulting error term requires a uniform gradient bound $|\nabla u_{k}|_{g}\leqslant C$. Obtaining this bound (without assuming the existence of a subsolution) seems to be a highly challenging problem in its own right. While this error term could be eliminated using the rectification theorem, this approach can not be applied directly since $v$ must have a zero on $X$. To address this issue, we perform the annulus trick as follows: let us consider a product manifold $\mathcal{X}:=X\times\mathbb{C}$. Let $\omega_{\mathbb{C}}:=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}|\tau|^{2}$ be the standard Kähler form on $\mathbb{C}$ and $w:=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$ a holomorphic vector field on $\mathbb{C}$ whose real part generates the action of the real torus $T^{\prime}\simeq\mathbb{S}^{1}$ on $\mathbb{C}$ by rotations. Set
@@
Since the holomorphic vector field $v+w$ has no zeros on $\mathcal{X}_{3}$, for any fixed $p\in\mathcal{X}_{3}$ we can choose local homomorphic coordinates $(w^{1},\ldots,w^{n+1})$ centered at $p$ such that $\frac{\partial}{\partial w^{1}}=v+w$ by the rectification theorem. In this coordinates, after adjusting the affine term, the local potential function $\phi$ of $\omega_{\mathcal{X}}$ is expressed as $\phi(w)=\overline{w}^{T}Gw+O(|w|^{3})$ for some positive definite Hermitian matrix $G=(G_{i\bar{j}})$. In particular, we have $|v+w|_{g_{\mathcal{X}}}^{2}(p)=G_{1\bar{1}}$, where $g_{\mathcal{X}}$ denotes the Riemannian metric with respect to $\omega_{\mathcal{X}}$. Take $B:=G^{1/2}$ so that the local potential function $\phi$ in the coordinates $z=Bw$ becomes $\phi(z)=|z|^{2}+O(|z|^{3})$. The holomorphic vector field $v+w$ still has constant coefficients in $(z^{1},\ldots,z^{n+1})$, which we denote by $c(v+w)\in\mathbb{C}^{n+1}$. Also a simple computation shows that $|c(v+w)|=|v+w|_{g_{\mathcal{X}}}(p)\leqslant\sup_{\mathcal{X}_{3}}|v+w|_{g_{\mathcal{X}}}$.
@@
For any $\varepsilon\in(0,\varepsilon_{7})$, we have
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\leqslant F_{\omega,-\varepsilon}(\chi_{u_{k}})+\varepsilon^{2}$
on $B_{j,4R}$ for all $j\in\mathcal{J}$, where the quantity in the left-hand side is calculated on $\mathcal{X}$.